# GPTVoiceTasker: LLM-Powered Virtual Assistant for Smartphone

### v1

@@ 3. The GptVoiceTasker System
We introduce GptVoiceTasker, a virtual assistant that empowers users to efficiently perform various tasks on their smartphones using voice commands. By applying different prompt engineering techniques (Section 3.1), GptVoiceTasker harness the power of LLMs to perform different logical tasks. Upon receiving a user command, GptVoiceTasker first attempts to automatically execute the task using the saved database (Section 3.3). Should a task prove unfeasible using the pre-existing database, GptVoiceTasker will perform a series of step-by-step predictions of on-screen navigation to complete a task (Section 3.2). Simultaneously, the system records these interactions for subsequent automated execution.
3.1. Prompt Engineering
@@
In this figure, we show an example of our prompt to LLMs to determine the most relevant target to press. The prompt includes a tasks description, the user command, the application information, the UI elements on current screen, an exemplar that includes a prompt and response for LLMs to follow. For the response format, Large Language Model return the chain of thought and the result, which include the information of the target such as the textual content and the bounds of UI element on the screen to perform a press.
While a naive approach to prompting LLMs for various tasks may yield suboptimal results due to low accuracy and randomness in responses (Chen et al., 2022), we propose the adoption of different prompt engineering techniques. These techniques involve crafting prompts according to specific rules and components to elicit optimal responses from LLMs (Liu and Chilton, 2022). We created multiple prompt templates [^2] https://github.com/vuminhduc796/GPTVoiceTasker/blob/main/prompts.txt, which is applied in both On-Screen Interaction (Section 3.2) and Usage-based Execution (Section 3.3). In Fig. 2, we illustrate an example of our prompt designed to determine the most appropriate target for tapping. This section explains different prompt engineering techniques applied in GptVoiceTasker, showcasing their significance in harnessing LLMs for complex reasoning tasks across multiple components in our system.
3.1.1. Least-to-most Prompting.
@@
In addressing complex tasks, such as predicting the action to perform on the screen from the user command, GptVoiceTasker employs the Least-to-most Prompting strategy (Zhou et al., 2022). This technique involves breaking down a complex task into smaller, manageable prompts. The goal is to enhance the model’s understanding of the logical flow, thereby improving accuracy. For example, GptVoiceTasker implements a two-step prompting approach for On-Screen Interactions (Section 3.2). First, the model is prompted to map user intent to a specific action (e.g., tapping an element, entering text, scrolling). Subsequently, based on the determined action, subsequent prompts are sent to identify the target UI element for executing that action (Fig. 2). This approach allows the LLMs to concentrate on atomic tasks, contributing to improved accuracy compared to a naive approach of identifying both action and target in the same prompt (Wang et al., 2023a).
3.1.2. Few-shot Prompting.
Few-shot prompts (Brown et al., 2020), incorporating completed task examples into the prompt (as in Examplar 1 in Fig. 2) itself, empower LLMs like GPT to facilitate rapid comprehension of specific rules and guidelines, thereby markedly improving output accuracy and reliability. For instance, when discerning the action in a user command, we include exemplars with synonyms and homophones to enhance adaptability and address speech recognition variations. This integration of few-shot learning, coupled with the LLM’s adept vocabulary and grammar comprehension, boosts versatility across diverse scenarios. Importantly, it effectively rectifies terminology discrepancies, a common issue in voice assistants using automatic speech recognition (Barros and Goguely, 2021). This advancement represents a notable improvement over heuristic-based methods, especially in navigating linguistic intricacies.
3.1.3. Chain of Thought.
When dealing with complex prompts that demand logical reasoning, LLMs often struggle to fully interpret and deliver accurate output. To address this, we integrate “Chain of Thought” (Wei et al., 2022) into our few-shot exemplars to help LLMs simulate human-like reasoning and provide logical output. By employing this mechanism, we ensure consistent and coherent reasoning, enabling a seamless progression of ideas and actions in generated responses. In GptVoiceTasker, the chain of thought is added to each response in examples where few-shot prompting is applied. In few-shot exemplar 1, as depicted in Fig. 2, we provided a chain of thought to predict the most relevant UI element to find the “video history in YouTube”. Despite the task requiring multiple steps to navigate to the history page, we asked the LLM to identify key indicators and execute logical reasoning specific to the task. This enabled it to accurately predict the most likely button for interaction. As a result, the system can repeat the same command, incorporate it into the prompts, proceed with the predicted actions, and continue until it reaches the intended destination.
3.2. On-screen Interaction
Upon receiving a task from users, GptVoiceTasker will automatically perform multiple on-screen action executions until the task is accomplished. We first collect run-time UI elements, execute reasoning tasks from user commands, and perform actions on the user’s device. GptVoiceTasker continues to repeat this iterative process to perform each action to accomplish the user tasks.
3.2.1. UI Semantic Extraction
To ensure a comprehensive understanding of the current UI elements and their semantics, GptVoiceTasker utilises Android Accessibility Service to continuously capture UI and analyzes new changes in the mobile screen, following previous research (Vu et al., 2023). Note that the raw UI elements we extracted may be afflicted by UI noise, which is a prevalent issue linked to the real-time gathering of UI elements (Li et al., 2022). This problem arises when the collected UI information does not align with its visual representation, which affects the semantic understanding of LLMs on current UI elements, resulting in incorrect interactions. To address this, we implement heuristics to mitigate potential inaccuracies and ensure the reliability of collected UI information. First, we utilise the collected coordination of each UI elements to eliminate out-of-bound or empty elements. We also eliminate views are fully overlapped by other views, which does are invisible and not interactible. In addition, we remove those views that do not contain any interprettable information, such as empty view containers.
Our primary emphasis is on representing mobile UI screens in a textual format that can be interpreted by Large Language Models (LLMs) through text-based input. While recent research proposed the translation of XML representation of UI screens to HTML format (Wang et al., 2023a; Feng and Chen, 2023) to overcome prompt length limitations, this approach becomes less relevant as modern models relax restrictions on prompt length, allowing developers to present more information in a prompt. For textual components like text fields and buttons, we gather essential semantic information from their labels and contents. We collect alternative text and resource names for icons and images, which offers meaningful information as defined by developers. Additionally, we augment the collected information with attributes such as the precise element location on the screen and a more comprehensive view hierarchical structure, previously absent in HTML representations. These details play a crucial role in illustrating relationships between UI elements, as demonstrated by the example of delivery times for each restaurant in the UberEats app (Fig. 3). This capability proves valuable in accommodating user commands, such as “Select the restaurant with the fastest delivery time”, a task unattainable with a flattened UI representation approach. Furthermore, we can cater for commands that user references to UI elements by their locations, such as “Press the icon at the top-right corner”. Lastly, each UI element is enriched with supported interaction types, indicating potential actions like PRESS or ENTER_TEXT to support LLM propose the appropriate action.
To enhance the contextual understanding and reasoning capabilities of the LLM, we not only fetch on-screen data but also retrieve relevant system-related information, including the app name and activity name of the current screen. By incorporating these system-related details, GptVoiceTasker gains a comprehensive understanding of the user’s current context, which further aids in validating each step in multi-step executions.
@@ 3.2.2. Action Executor
We combine the user command with UI and contextual information collected in previous step and applies different prompt engineering techniques as described in Section 3.1 to create prompts. These prompts are then sent to LLMs, which will first detect most appropriate actions and then detects the associated target UI element for this action. GptVoiceTasker uses this information to perform the interactions on user’s device, such as tapping, scrolling, or entering text on certain UI element. Prior mobile automation approaches (Pan et al., 2022) encounter challenges in handling runtime UI changes and app updates that might alter UI representations and operation sequences. In contrast, our approach exhibits robustness in the face of such issues, as the sequence of steps is dynamically dependent on the UI on the screen. This adaptability allows GptVoiceTasker to navigate through dynamic UI changes and updates, ensuring reliability in task automation on smartphones. Additionally, GptVoiceTasker provides audio feedback to users, confirming that the system is automatically proceeding to the next command. This real-time feedback ensures a smooth and intuitive user experience with GptVoiceTasker’s interaction capabilities.
@@
To address challenges inherent in automation within real-time systems, including the detection of failures in each automated step (Xiao and Wang, 2008), our implementation integrates a screen transition detector. This detector utilizes the Hamming distance (Kopelowitz and Porat, 2018) to gauge the difference between two screens, ensuring that the UI accurately reflects changes resulting from the automated action. Moreover, for validating the success of actions, we implement additional heuristics. For example, in scenarios involving ENTER_TEXT, we verify the presence of the entered text in the target text box. If an action proves inexecutable (i.e., not inducing appropriate changes to the UI), we iteratively repeat the step with supplementary information about the failed action, thereby excluding these actions from the selection.
Another challenge in mobile app automation arises from the prevalent dependency of mobile UI screens on live internet content, often loaded asynchronously after the screen appears on smartphones. The collection and utilization of incomplete or loading UIs can compromise the accuracy of detecting suitable actions. In response, GptVoiceTasker introduces an innovative approach by delaying subsequent actions until the screen is fully loaded. This is achieved by detecting screen-loading widgets in Android, leveraging information such as widget names, types, and shapes to detect progress bars and loading indicators. Furthermore, we enhance our approach by collecting data on network transmissions to identify ongoing content downloading tasks, inspired by the methodology presented in (Liu et al., 2020), which leverages network analysis for detecting ads. These methods significantly enhance the reliability of GptVoiceTasker in effectively navigating through the app’s interface.
3.3. Usage-based Execution
@@
In mobile apps, the UI elements on a specific app page are predefined by developers when developing an app page. Therefore, the series of user interactions on the screen to achieve a task will be consistent across different times. Based on this matter, GptVoiceTasker automatically create a saved path for each user command, allowing GptVoiceTasker to replicate interactions when receiving a similar command from users. Unlike previous approaches (Pan et al., 2022; Li et al., 2017) that depend on a manual task creation process for automation support, GptVoiceTasker automatically records app transitions through on-screen navigation in Section 3.2. This automated process not only allow wider coverage of automated tasks but also eliminates the need for manual efforts in predefining shortcut tasks. In this section, we outline our method (as in Fig. 3) to streamline subsequent similar tasks from users, which combines both LLMs-based and heuristic-based modules. We introduce our database design in Section 3.3.1. First, we identify the current UI screen displays on user device and the destination screen (Section 3.3.2). After that, we find the most viable path from current screen to the destination screen (Section 3.3.3). Finally, we use incorporate human interactive feedback to validate and finetune the execution for further usages (Section 3.3.4).
Figure 3. An example use case in Uber Eats to how GptVoiceTasker use the historical tasks to execute user new command.
@@
Figure 3. An example use case in Uber Eats to how GptVoiceTasker use the historical tasks to execute user new command.
@@ 3.3.1. Transition Graph
In line with previous approaches to automating tasks in mobile apps (Chen et al., 2019), GptVoiceTasker incorporates a directed graph for each app to facilitate seamless transitions between different app pages. Each node in the database corresponds to a UI page in the app, containing a unique ID and a screen description generated as outlined in Section 3.3.2. Moreover, for each node, we maintain a list of previously used commands by users that concluded on this page. This compilation signifies that the requested functionality from these user commands is associated with this specific app page. The directed edges between nodes represent potential transitions from one UI page to another, capturing information about the associated action type and the target UI element for the transition. These edge details empower GptVoiceTasker to replicate user actions and screen traversals. As users engage with the app, this graph dynamically expands, automatically incorporating new UI pages and associated transitions.
@@ 3.3.2. Screen Description & Command Pattern Matching
While previous automation approaches focus on executing tasks starting from the launcher page of the app (Pan et al., 2022; Arsan et al., 2021), they may not be suitable for cases where users are currently on other pages, necessitating additional navigation steps to return to the launcher page. To address this challenge, we first identify the current screen of the app that the user is viewing in our graph database. However, heuristically comparing and identifying a UI page using the on-screen textual information collected from the XML representation proves impractical due to the contextual dynamicity of information within UI elements on a specific app page. For example, consider the search result page in the Uber Eats app, where distinct restaurants are displayed for “spaghetti” and “sushi” search results. Despite featuring different content, these pages share the same layout and, therefore, should be identified as the same node in our graph database. To achieve this, we leverage the capabilities of LLMs to semantically summarize the UI content (Wang et al., 2023a) into a semi-structured description. We collect content from the XML hierarchical representation of the screen, along with additional context about the current activity and app name, to generate a high-level description of the overall functionality that the screen serves. Subsequently, GptVoiceTasker summarizes the list of interactive elements, including clickable, scrollable, and text-editable elements, and appends them to the description. To locate the current UI page within the list of visited screens in an app, we prompt LLMs to perform semantic matching between the current screen description and the saved list of screen descriptions from the app. This process allows us to trace back the ID of the screen in the graph database, effectively addressing the navigation challenges from any point within the app.
After identifying the current screen, we perform semantic matching to find destination screen. We define destination screen as the most relevant screen that can serve user request, for example, the destination screen for user command “I want to order spaghetti” is the search result page for the keyword “spaghetti”. We utilise the user saved commands and screen description for each app page to prompt LLMs to determine the most relevant app page from the app. We build a prompt that includes user request and the list of saved screen in the app to rank the relevancy of each app screen with the request that user made.
@@
After identifying the current and destination pages within the saved graph database, GptVoiceTasker utilizes the Shortest Path algorithm (Golden, 1976) to determine the sequence of actions required to navigate from the start to the destination node. GptVoiceTasker extracts this sequence from the edges connecting the start node to the destination node and executes each action in sequence using the Action Executor described in Section 3.2.2.
Given the dynamic nature of smartphone GUIs, where the relative coordinates of buttons may vary, especially in scrollable screens or due to unexpected pop-ups and ads, execution validation is crucial. To address this, GptVoiceTasker performs validation on each UI page in the path, ensuring the automation follows the expected sequence. After each action in the sequence, the screen description is generated based on current displayed UI (as explained in Section 3.3.2) and compared to the expected description stored in the graph node representing the anticipated UI. If the descriptions do not match, the tool navigates back to the previous app page and prompts LLMs to propose an action. The prompt created for on-screen action predictions is then used to feed the action to the LLM, regenerating the appropriate action. This iterative process ensures that the automation adapts to changes in the smartphone GUI, allowing GptVoiceTasker to dynamically execute actions based on real-time screen information.
Additionally, GptVoiceTasker leverages command parameterization techniques (Pan et al., 2022; Krosnick and Oney, 2022), allowing saved paths to be reused for similar tasks with different parameters. Utilising the enriched vocabulary and robust natural language understanding capabilities of the LLM, GptVoiceTasker prompts the LLM to function as an advanced Named Entity Recognition (NER) system. This system identifies and replaces all occurrences of substitutable words with new keywords in the action sequence. For example, in the Uber Eats app scenario shown in Fig. 3, where a user has a saved command for “I want to order some sushi”, GptVoiceTasker recognizes “sushi” as a parameter value representing the type of food, which is replaced by “spaghetti”, transforming the command from ENTER “sushi” to ENTER “spaghetti”. Subsequently, GptVoiceTasker executes this adapted sequence of actions to accomplish the task.

### v2

@@ 3. The GptVoiceTasker System
Figure 2. An example use case in Home Workout application when the user needs to interact with the smartphone hands-free due to physical busyness. When performing an unprecedented tasks (Section 3.1), GptVoiceTasker repeatedly predicts on-screen actions with current UI information and executes the response to achieve user tasks. The interactions collected during this process is then saved to streamline the execution of subsequent similar tasks (Section 3.2).
[…]
@@
We introduce GptVoiceTasker, a virtual assistant that empowers users to efficiently perform multi-step tasks on their smartphones using voice commands. Upon receiving a user command, GptVoiceTasker first attempts to streamline the task using the collected in-app navigation database (Section 3.2), to improve execution efficiency and reliability. If a task is unprecedented and not in our saved records, GptVoiceTasker will perform a series of step-by-step predictions of the on-screen navigation sequence to complete the task (Section 3.1). Simultaneously, the system expands the database with new in-app navigation knowledge for subsequent autonomous task execution.
3.1. Unprecedented Task Exploration
Upon receiving an unprecedented task from the user, GptVoiceTasker will progressively predict and automatically execute each step until the task is accomplished. For each step, we collect contextual data from the mobile UI, task execution context, and current application information, which is combined with system-level information. GptVoiceTasker constructs all relevant data into prompts in a specific format and feeds them to the LLMs to determine the appropriate action on the user’s smartphone. Upon receiving the response from the LLMs, GptVoiceTasker executes the action on the smartphone accordingly.
3.1.1. Data Collection Module
As LLMs execute logical reasoning based on textual inputs, known as prompts, a detailed and comprehensive prompt aids LLMs in understanding the task at hand and generating appropriate responses. Therefore, our primary focus is to incorporate sufficient information into our prompts to ensure accurate decisions from LLMs. This critical information, which we define as knowledge, is categorized into User Interface (UI) knowledge, Task knowledge, Application knowledge, and System knowledge. We extract this information through static analysis of the smartphone and its applications.
UI knowledge. Information about the on-screen UI elements is a major component in our prompts, as it allows LLMs to comprehend the content currently displayed. Our primary emphasis is on representing smartphone GUIs in a textual format that can be interpreted by LLMs through text-based input. While recent research has proposed converting the UI elements list to HTML format (Wang et al., 2023a; Feng and Chen, 2023) to reduce the prompt length, this approach becomes less relevant as LLMs now have relaxed restrictions on the number of tokens in a prompt. Therefore, we propose a more comprehensive view of the hierarchical structure of smartphone GUIs to improve the decision accuracy of LLMs. We represent each screen as a tree of nodes, with non-leaf nodes representing UI containers and leaf nodes representing visible UI elements. For each UI element, we collect the element type, text label, and append it with a unique ID. For some element types, such as buttons or text fields, the label can be extracted directly from the screen. However, for certain graphical UI elements like icons or image buttons, such information is not readily available. A potential solution is to apply deep learning models to predict the potential label of icons (Chen et al., 2022b). Nevertheless, this approach can cause excessive overhead on app pages that include multiple images and icons, significantly impacting the responsiveness of real-time assistants. Alternatively, we propose a lightweight approach to collect alternative captions and resource names of these elements as labels, as they often include informative descriptions. For example, the search icon in screen 1 of Figure. 2 has the resource name “ic_search”, defined by app developers, indicating that this button is used for the search functionality. Furthermore, we collect the precise element location on the screen to cater to commands that refer to UI elements by their locations, such as “Press the icon at the top-right corner”. We also retrieve the list of allowed actions for each element, which can include CLICKABLE, TEXT_EDITABLE, SCROLLABLE, etc. This knowledge acts as a guardrail to ensure that LLMs do not return unsupported actions, such as pressing a disabled button. The runtime UI elements may contain UI noise, which is a prevalent issue linked to the real-time gathering of UI elements (Li et al., 2022). This problem arises when the collected UI information does not align with its visual representation, which affects the semantic understanding of LLMs on current UI elements, resulting in incorrect interactions. To address this, we implement heuristics to mitigate potential inaccuracies and ensure the reliability of collected UI information. First, we utilise the collected coordination of each UI element to eliminate out-of-bound or empty elements. We also eliminate views are fully overlapped by other views, which does are invisible and not interactible. In addition, we remove those views that do not contain any interpretable information, such as empty view containers.
Task knowledge. As complex tasks on smartphones involve multiple steps, treating each step as a separate action may lead to execution inaccuracies due to the misalignment of sequential actions. Additionally, the system may reattempt a single action multiple times, potentially causing an endless execution loop. To address this, we maintain information about the currently executing task and include it in our prompts. The task knowledge in each prompt specifies the user’s request, the previously executed actions, and the visited pages. Such knowledge is incorporated into our prompts as natural language in a predefined template (e.g., “Started from ¡page 1¿, we ¡action 1¿ ¡target 1¿ to get to ¡page 2¿. After that, we ¡action 2¿ ¡target 2¿ to get to ¡page 3¿.”), helping LLMs to comprehend the previous actions and better align the subsequent actions accordingly.
Application knowledge. Our experiment shows that LLMs can incorporate the knowledge about smartphone apps to output better results. This is particularly helpful for popular apps as LLMs are more familiar with these apps in their training data. In addition, these information helps validating each step in multi-step executions. For example, if the action caused unintended navigation to another app, LLMs can utilise the current app name to know this changes and output appropriate action to navigate back to the requested app. Therefore, we provide the app name, package name and list of activities as the application knowledge in our prompt.
System knowledge. To enhance the contextual understanding and reasoning capabilities of the LLMs, we not only fetch on-screen data but also retrieve relevant system-related information. We collect resolution information, which identifies the current screen orientation, and include it in our prompts. In addition, the screen resolution helps in the pre-processing of collected UI elements, such as validating the position of UI elements for out-of-bounds checking.
3.1.2. Private Information Anonymisation.
Smartphones often contain and display personal data, including phone numbers, addresses, and payment details. Anonymizing this data is crucial when inputting UI screen content into LLMs to prevent exposing sensitive user information. To this end, we have integrated a lightweight Named Entity Recognition (NER) to efficiently searches and categorizes textual information within the UI elements. Upon detecting private information, we substitute it with standardized tags, such as ¡address¿ or ¡phone number¿, before the data is processed by the LLMs. This approach not only preserves user privacy but also ensures that LLMs understand the UI semantics accurately to generate appropriate responses.
3.1.3. Prompt Creation
@@
In this figure, we show an example of our prompt to LLMs to determine the most relevant target to press. The prompt includes a task description, the user command, the application information, the UI elements on the current screen, and an exemplar that includes a prompt and response for LLMs to follow. For the response format, Large Language Model return the chain of thought and the result, which includes the information of the target such as the textual content and the bounds of UI element on the screen to perform a press.
@@
We include the user command with collected data in the previous step in our prompt to predict the most suitable action to perform on user smartphone. As the naive approach to prompt LLMs for various tasks may yield suboptimal results due to low accuracy and randomness in responses (Chen et al., 2022a), we adopt of different prompt engineering techniques. These techniques involve crafting prompts according to specific rules and components to elicit optimal responses from LLMs (Liu and Chilton, 2022). Following Least-to-most Prompting strategy (Zhou et al., 2022), GptVoiceTasker implements a two-step prompting approach to determine the action. First, we map the user task to a specific action (e.g., tapping an element, entering text, scrolling). Subsequently, based on the determined action, subsequent prompts are sent to identify the target UI element for executing that action ( as illustrated as an example Figure. 3). This approach allows breaking down a complex task into smaller steps, enabling the LLMs to improve accuracy.
To empower LLMs to facilitate rapid comprehension of specific rules and guidelines, we employed the Few-shot Prompting approach (Brown et al., 2020), where we provide additional examplars in our prompt. Each examplar contains the sample prompt as the input with its corresponding response as the expected output. In addition, we integrate “Chain of Thought” (Wei et al., 2022) into our few-shot exemplars to help LLMs simulate human-like reasoning and provide logical output. This includes a sentence that explains the logic behind its output, which guides the LLM to apply a similar thought process in handling tasks. We dynamically calculate the number of examplars in a prompt based on the estimation of tokens used to include the knowledge. This approach maximizes the number of exemplars, thereby enhancing the accuracy of the response, while ensuring we do not exceed the token limit. In few-shot exemplar 1, as in Figure. 3, we provided a chain of thought that explains why Library button should be pressed to find the video history in YouTube.
@@ 3.1.4. Action Executor
GptVoiceTasker extracts the action and target from LLMs response to perform the interactions on user’s device, such as tapping, scrolling, or entering text on certain UI element. Prior mobile automation approaches (Pan et al., 2022) encounter challenges in handling runtime UI changes and app updates that might alter UI representations and operation sequences. To improve from previous approaches, GptVoiceTasker dynamically propose the action based on the current UI, ensuring reliable navigation through dynamic UI changes and updates. Additionally, GptVoiceTasker provides audio feedback to users, confirming that the system is automatically proceeding to the next command. This real-time feedback ensures a smooth and intuitive user experience with GptVoiceTasker’s interaction capabilities.
@@
To address challenges in real-time system automation, including failure detection in automated steps (Xiao and Wang, 2008), our implementation integrates a screen transition detector. This detector employs Hamming distance (Kopelowitz and Porat, 2018) to measure differences between screens, ensuring screen content changes due to the action. Moreover, to validate the success of actions, we implement additional heuristics. For example, in scenarios involving ENTER_TEXT, we verify the presence of the entered text in the target text box. If an action proves inexecutable (i.e., not inducing appropriate changes to the UI), we repeat the step with supplementary information about the failed interaction attempts, thereby excluding these actions from the selection. Another challenge in mobile app automation is the dependency of UI screens on asynchronous internet content. Providing LLMs with incomplete or loading UI instead of fully rendered screen content may reduce the precision in identifying appropriate actions. In response, GptVoiceTasker delays UI collection until the screen is fully loaded. This is achieved by detecting screen-loading widgets in Android, leveraging information such as widget names, types, and shapes to identify progress bars and loading indicators. Additionally, we enhance our approach by collecting data on network transmissions to identify ongoing content download tasks, inspired by the methodology presented in (Liu et al., 2020), which leveraged network analysis for detecting ads. These methods enhance the reliability of GptVoiceTasker in effectively navigating through the app’s interface.
3.2. Precedented Task Automation
@@
In mobile apps, UI elements on a specific app page and in-app navigation paths are predefined by developers during app development. Therefore, the series of user interactions on the screen to complete a task in one app remains consistent over different instances. Based on this, GptVoiceTasker automatically creates a saved path for each user request, enabling GptVoiceTasker to replicate interactions when receiving the same or similar commands from users. Unlike previous approaches (Pan et al., 2022; Li et al., 2017), which depend on manual task creation for automation support, GptVoiceTasker automatically records in-app navigation through on-screen interactions as described in Section 3.1. This automated process not only expands the coverage of automated tasks but also eliminates the need for manual efforts in pre-defining shortcut tasks. In this section, we outline our method (as shown in Figure. 4) to streamline subsequent similar tasks from users, combining both LLMs-based and heuristic-based modules. We introduce our database design in Section 3.2.1. Initially, we identify the current UI screen displayed on the user’s device and the destination screen (Section 3.2.2). We then find the most viable path from the current screen to the destination screen (Section 3.2.3). Finally, we incorporate human interactive feedback to validate and fine-tune the execution for future use (Section 3.2.4).
Figure 4. An example use case in Uber Eats to how GptVoiceTasker use the historical tasks to execute user new command. The system first locate the current screen and destination screen from the collected graph. After that, it identifies and execute the action sequence to traverse to the destination screen. Finally, we utilise feedback from users to improve subsequent execution.
[…]
@@ 3.2.1. Transition Graph
In alignment with previous approaches for automating tasks in mobile apps (Chen et al., 2019), GptVoiceTasker utilizes a directed graph for each app to enable seamless transitions between different app pages. Each node in the database corresponds to a UI page in the app, containing a unique ID and a screen description, as detailed in Section 3.2.2. Additionally, for each node, we record the list of previous user requests that concluded on this page, as these indicate the functions served by the page. The directed edges between nodes signify possible navigation paths from one UI page to another. These edges hold details about the required actions and target UI elements for page transitions. This data enables GptVoiceTasker to perform specific actions on the identified targets, thus replicating user actions to facilitate navigation between screens. This graph expands automatically to include new UI pages and associated page transitions as users interact with their smartphones.
@@ 3.2.2. Screen Description & Command Pattern Matching
Previous task automation approaches typically start from an application’s launcher page (Pan et al., 2022; Arsan et al., 2021), a method that falls short in real-world scenarios, especially when users request voice assistants to complete ongoing tasks. To address this, GptVoiceTasker enables task automation from any page of the app, aiming to fulfill any new or ongoing tasks from users. We first identify the node that represent the current app page in our graph database. However, identifying a page by performing string matching on the UI content presents substantial practical challenges. This difficulty arises as a specific application page can display dynamic content. For example, consider the search result page in the Uber Eats app, where distinct restaurants are displayed for “spaghetti” and “sushi” search terms. Despite featuring different UI content as they display different restaurants, these pages share the same layout and are identified as the same node in our graph database. To compare between app pages, GptVoiceTasker leverages the capabilities of LLMs to semantically summarize the UI content into a semi-structured description in natural language. The description includes a short paragraph describing overall functionality that the screen serves based on the UI elements, current activity name and app name. In addition, GptVoiceTasker appends the list of interactive elements, including clickable, scrollable, and text-editable elements to the description. To locate the current UI page among previously visited pages, GptVoiceTasker uses LLMs for semantic matching between the screen’s description and existing screen descriptions in the database. As LLMs are proficient in logical tasks on natural language, the response allows us to identify the node in the graph database.
After identifying the current screen, we perform semantic matching to find the destination screen. We define the destination screen as the most relevant screen that can serve user requests. For example, the destination screen for user command “I want to order spaghetti” is the search result page for the keyword “spaghetti”, where users can view the list of spaghetti restaurants. We utilise the saved commands and screen description for each app page to prompt LLMs to determine the most relevant page from the app. We build a prompt that includes user requests and the list of saved screens in the app to rank the relevancy of each app screen with the request that user made.
@@
After identifying the current and destination pages within the saved graph database, GptVoiceTasker uses the Shortest Path algorithm (Golden, 1976) to determine the sequence of actions required to navigate from the start to the destination node. GptVoiceTasker extracts this sequence from the edges connecting the start node to the destination node and executes each action in sequence using the Action Executor described in Section 3.1.4.
Given the dynamic nature of smartphone GUIs, where the relative coordinates of buttons may vary due to scrollable screens or unexpected pop-ups and ads, execution validation is crucial. To ensure automation follows the expected sequence, GptVoiceTasker performs validation after each action. We generate the description for the current page (as explained in Section 3.2.2) and compare it with the description in the anticipated node. If the two descriptions do not semantically match, GptVoiceTasker reverts the preceding action and seeks an alternative path to the destination page. If no other path is found, GptVoiceTasker predicts and executes subsequent steps using the On-screen Interaction Module to explore a new path. This iterative process adapts to changes in the smartphone GUI, dynamically executing actions based on real-time screen information.
Additionally, GptVoiceTasker leverages command parameterization techniques (Pan et al., 2022; Krosnick and Oney, 2022), allowing saved paths to be reused for similar tasks with different parameters. Using the LLMs’ enriched vocabulary and robust natural language understanding, GptVoiceTasker prompts the LLMs to function as an advanced Named Entity Recognition system. This system identifies and replaces substitutable words with new keywords in the action sequence. For instance, in the Uber Eats app scenario (Figure. 4), a saved command for I want to order some sushi” can be adapted by replacing sushi” with spaghetti”, thus modifying the command from ENTER sushi” to ENTER “spaghetti”. GptVoiceTasker then executes this adapted sequence to complete the task.